OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 20 and 2500, respectively.

OpenStudy (anonymous):

2500/20 = 125 125 = 5^3 an fifth term / second term = r^3 where r is the common ratio so r^3 =5^3 hence r= 5 and the first term is 20/5=4 so nth term = 4 * 5^(n-1)

OpenStudy (anonymous):

ahhh okay. That makes sense :) thanks

OpenStudy (anonymous):

yw